# A backdoor attack against link prediction tasks with graph neural networks

### v1

@@ Abstract
Graph Neural Networks (GNNs) are a class of deep learning models capable of processing graph-structured data, and they have demonstrated significant performance in a variety of real-world applications. Recent studies have found that GNN models are vulnerable to backdoor attacks. When specific patterns (called backdoor triggers, e.g., subgraphs, nodes, etc.) appear in the input data, the backdoor embedded in the GNN models is activated, which misclassifies the input data into the target class label specified by the attacker, whereas when there are no backdoor triggers in the input, the backdoor embedded in the GNN models is not activated, and the models work normally. Backdoor attacks usually occur when the training process of GNN models is not controlled, for example, backdoor attacks can occur when using models trained by third parties or using third-party data to train models. Backdoor attacks are highly stealthy and expose GNN models to serious security risks. Currently, research on backdoor attacks against GNNs mainly focus on tasks such as graph classification and node classification, and backdoor attacks against link prediction tasks are rarely studied. In this paper, we propose a backdoor attack against the link prediction tasks based on GNNs and reveal the existence of such security vulnerability in GNN models, which make the backdoored GNN models to incorrectly predict unlinked two nodes as having a link relationship when a trigger appear. The method uses a single node as the trigger and poison selected node pairs in the training graph, and then the backdoor will be embedded in the GNN models through the training process. In the inference stage, the backdoor in the GNN models can be activated by simply linking the trigger node to the two end nodes of the unlinked node pairs in the input data, causing the GNN models to produce incorrect link prediction results for the target node pairs. We demonstrate that this attack is effective and efficient by performing experimental evaluations on four popular models and four public datasets and comparing them with baseline, and the experimental results show that in the black-box scenario, this attack can achieve high attack success rates (more than 89 %) with small model accuracy losses (less than 1%) and small poisoning rates (about 1%).
1 Introduction
@@
Graph-structured data is ubiquitous in the real world, covering various application domains such as social networks [1],urban transportation networks [2], knowledge graphs [3], and biological networks [4]. In order to efficiently process graph-structured data, various powerful graph neural network(GNN) models have been proposed. GNNs adopt the message passing mechanism to update the representations of nodes by aggregating information from their neighbors. GNNs can be used for various graph related tasks such as graph classification [5], node classification [6], link prediction [7, 8] etc. Link prediction aims to predict the existence of links between pairs of nodes in a graph, which is an important task of GNNs and has a wide range of applications such as friends recommendation in social networks [9], prediction of missing triples in knowledge graphs [10], and prediction of protein functions as well as interactions in biological networks [11].
@@
GNNs have powerful performance in processing graph-structured data. However, recent studies [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22] have shown that GNNs are vulnerable to a new security threat called backdoor attacks. In backdoor attacks, the attacker injects specific patterns (called backdoor triggers) in the training data and then embeds the backdoor in the GNN models through training. In the inference phase, when the input samples contain the backdoor trigger, the backdoor in the models will be activated and perform the action specified by the attacker in advance. For example, the GNN models embedded with the backdoor misclassify input samples with triggers as attacker-specified target class label, whereas they perform normally when processing samples that do not contain backdoor triggers. Backdoor attacks often occur when the training process of GNN models are uncontrolled, e.g. backdoor attacks can occur by using models trained by third parties or by using third-party data to train the models [22]. Backdoor attacks are highly stealthy and they have posed serious security threat to GNNs.
@@
To the best of our knowledge, the only existed backdoor attack studies on link prediction are LB [20] and DLB [21], where DLB is a backdoor attack method for link prediction on dynamic graphs, and LB is a backdoor attack method for link prediction on static graphs, which studies a similar problem to that of our work. Specifically, LB optimizes the structure and features of a random initial subgraph by using the gradient information of the target model, then obtains a dense subgraph as the trigger for the backdoor attack, and finally poisons the graph dataset with this subgraph, so as to embed the backdoor into the GNN models through training. Although LB achieves good results, this method requires a large attack budget because LB uses the gradient information of the target model to optimize the subgraph, and the dense subgraph makes the trigger easy to be detected.
@@
In this paper, we investigate whether there are less costly and more stealthy method to implement backdoor attacks against link prediction. Specifically, we propose a backdoor attack method for link prediction with a single node as the trigger, which targets backdoor models that incorrectly predict link relationships between two unlinked nodes as linked when the trigger occurs. Initially, a new node is created as the trigger, and the features of the trigger node are generated by counting the frequency of occurrence of each feature in the original graph and selecting the subset of features with fewer occurrences; Then appropriate node pairs are selected in the training graph to inject that trigger for poisoning, and the backdoor is embedded in the model through the training process. In the inference phase, the implementation of the backdoor attack can activate the backdoor in the model by simply connecting the trigger node to the two end nodes of the target node pairs in the test data, causing the model to produce incorrect prediction results for that target node pairs. The illustration of our backdoor attack is shown in Figure 1. The left side of the figure shows the input graph and the red node is the backdoor trigger node. Part (a) of Figure 1 represents the input graph without the trigger, and node A and node B form a target node pair between them; Part (b) of Figure 1 represents the input graph with the trigger, where the red trigger node connect node A and node B of the target node pair respectively. When predicting the link relationship between nodes A and B with the model embedded with the backdoor, the backdoor model works correctly and correctly predicts that there is no link between node A and node B because there is no trigger present, while for the input of (b), the backdoor in the backdoor model is triggered due to the presence of the trigger and the backdoor model incorrectly predicts a link exists between node A and node B. We conduct experimental evaluations on four popular models and four benchmark datasets and compare them with baseline to demonstrate that the attack is effective and efficient, and the experimental results show that the attack success rate can reach more than 92 % in the black-box scenario with the poisoning rates of about 1%, and that the backdoor models accuracies loss less than 1%.
@@
    We propose a backdoor attack against link prediction with GNNs, revealing that GNNs are vulnerable to this security risk. The backdoor attack has following three features: the attack achieves better stealthiness by using only one single node as the trigger, and it requires lower poisoning rate in training dataset ; The attack can be implemented in black box scenarios, where the attackers do not need to know the parameters of target GNN models or create surrogate models, so it is easier or more practical for attacker to launch the attack; The attack has lower computing cost without involving any computation of gradients.
@@
    We propose a simple and effcient method for contaminating training data, which can achieve better attack performance with lower poisoning rates by connecting the trigger node to the selected nodes pairs in the graph.
@@
The rest of the paper is organized as follows. In Section 2, we introduce the background knowledge of GNNs, link prediction and backdoor attacks; in Section 3, we introduce the adversarial attacks that GNNs are exposed to. Then we describe in detail the principle of backdoor attacks and the current state of study of backdoor attacks; in Section 4, we describe in detail our proposed method; in Section 5, we experimentally evaluate the performance of the proposed backdoor attack to demonstrate the effectiveness and feasibility of this backdoor attack; finally, Section 6 concludes the paper.
Figure 1: Illustration of our backdoor attack against link prediction. The trigger is the red node. Part (a) is the input graph without the trigger and node A and node B form a target node pair. The backdoor model works correctly and correctly predicts that there is no link between the node A and node B. Part (b) is the input graph obtained by adding the trigger from the graph in part (a) and the backdoor in the backdoor model is triggered due to the presence of the trigger and the backdoor model incorrectly predicts a link exists between the node A and node B.
2 Background
2.1 GNNs
Given an undirected and unweighted attribute graph $G=(V,E,X)$,where $V=\{v_{1},v_{2},\cdots,v_{N}\}$ is the set of nodes, $E$ is the set of edges, $X=R^{N\times d}$ represents the nodes attribute matrix, and the adjacency matrix $A\in R^{N\times N}$. For two nodes $v_{i},v_{j}\in V$, if $(v_{i},v_{j})\in E$, it means that there exists an edge between $v_{i}$ with $v_{j}$ and $A_{ij}=1$, otherwise, $A_{ij}=0$.$N=|V|$ is the total number of nodes, and $d$ is the number of node feature dimensions.
@@ 2.2 Link prediction
2.3 Link prediction methods based on GNNs
@@
$L(\theta)=\sum_{(i,j)\in E^{+}}{logp_{\theta}(A_{ij}=1|X)}+\sum_{(i,j)\in E^{-}}{log(1-p_{\theta}(A_{ij}=1|X))}$ (4)
@@
2.4 Backdoor attacks
Backdoor attacks are a type of poisoning attack against deep learning models, which aims to implant some hidden malicious functions, i.e., backdoors, in the target models by using poisoned samples with triggers during the models training process. In the inference phase, for input samples that contain trigger, the backdoor will be activated, causing the target models to produce incorrect prediction results, e.g., the input samples are incorrectly misclassified to the target class label specified by the attacker, whereas the backdoor model works normally for benign samples that do not contain trigger. Backdoor attacks were first proposed in the field of images [23, 24]. Gu et al. [23] proposed the backdoor attack for the first time in the image domain, where the backdoor model is trained by poisoning some training samples using special markers as triggers, and in the inference phase, the backdoor model classifies the images with the triggers as the target label, and classifies the clean images normally. Subsequently, backdoor attacks in the textual domain have also been extensively studied [25, 26], e.g., Qi.et.al [26] used syntactic structures as the triggers for textual backdoor attacks.
Backdoor attacks are highly stealthy and harmful because, in the absence of triggers, the performance and behavior of backdoor models are not significantly different from normal models and are not easily detected and defended against. Backdoor attacks can occur in many scenarios, such as when using datasets, platforms, or models provided by third parties for training or deployment. Backdoor attacks pose a serious threat to the security and trustworthiness of deep learning systems, and require sufficient attention and concern.
3 Related works
@@ 3.1 Adversarial attacks
3.2 Backdoor attacks
In the graph data field, backdoor attacks have been less studied, and current works focuse on backdoor attacks against graph classification [12, 13, 14, 15, 16] and node classification [16, 17, 18, 19]. Xi et al. [12] proposed for the first time a backdoor attack method for GNNs that uses subgraphs as the triggers that can be dynamically customized for different graphs with different triggers to poison the data. Yang et al. [13] revealed a transferable graph backdoor attack without a fixed trigger pattern, which implements black-box attacks on GNN by the attacking surrogate models. Zheng et al [14] rethink the triggers from the perspective of motifs (motifs are frequent and statistically significant subgraphs of graphs containing rich structural information), proposed motif-based backdoor attacks, and gave some in-depth explanations of backdoor attacks. Xu et al.[16] select the optimal trigger position through interpretable methods to accomplish backdoor attacks against graph classification and node classification.
The above studies mainly focus on graph classification and node classification, and the only current work on backdoor attacks under the context of link prediction are LB [22] and DLB [21]. DLB investigated the link prediction backdoor attack on dynamic graphs, it generates different initial triggers by Generative Adversarial Network (GAN) [37], and then based on the gradient information of the attack discriminator in the GAN, selects some of the links of the initial triggers to form a trigger set to accomplish the backdoor attack. LB is similar to our work, but LB uses subgraphs as triggers and generates triggers by using the gradient information of the target models. Compared to his work, our proposed backdoor attack specifies the following different features: the use of a single node as the trigger, the low poisoning rate of the training samples, and more unnoticeable; the ability to implement a black-box backdoor attack without the need for the specific parameters of the target GNN models or constructing surrogate models, which makes the attack easier; and the lack of the need to use the gradient information, which reduces the amount of computation and improves the efficiency of the attack.
4 Method
In this section, we explain in detail how our attack is realized. Table 1 summarizes the the notions used in the following sections and their explanations.
Table 1: The explanations of the notions
Notation Explanations
$G=(V,E,X)$ Graph $G$ with edge set $E$, node set $V$ and attribute matrix $X$
@@ 4 Method
$f_{c}$ The clean link prediction GNN model
@@
$x_{t}$ The features of the trigger node
$f_{b}$ The backdoored link prediction GNN model
$T$ Target link state
$E_{ij}$ The original-unlinked target node pair
$\widetilde{E_{ij}}$ The original-unlinked target node pair after embedding the trigger
$\alpha_{j}$ Frequency of occurrence of the i-th feature
$M_{k}$ the position indices of the nodes with the least frequent occurrence
@@
$NPS$ The Target node pair scoring function
4.1 Attack overview
The backdoor attack proposed in this paper aims to the link prediction task under transductive learning, the goal of which is to incorrectly predict a link relationship between two unlinked target nodes. Formally, given a graph $G$ and a trigger $n_{t}$, the attacker generates the poisoned graph $\tilde{G}$ by embedding the trigger $n_{t}$ into the graph $G$. The backdoor model $f_{b}$ is obtained by training on $\tilde{G}$ and the clean model can be denoted as $f_{c}$. The backdoor attack target can be denoted as:

### v2

# Effective backdoor attack on graph neural networks in link prediction tasks

@@ Abstract
Graph Neural Networks (GNNs) are a class of deep learning models capable of processing graph-structured data, and they have demonstrated significant performance in a variety of real-world applications. Recent studies have found that GNNs models are vulnerable to a new type of threat called backdoor attacks, where the adversary can inject hidden backdoor into the GNNs models so that the attacked model performs well on benign samples, whereas it performs pre-specified malicious behavior such as misclassification to adversary-specified target category if the hidden backdoor is activated by samples with the attacker-defined pattern (such as subgraphs) called backdoor trigger. Backdoor attacks usually occur when the training process of GNNs models is not full controlled by users, such as training on third-party datasets or adopting third-party GNNs models, which poses a new and serious threat to GNNs models.
@@
Currently, research on backdoor attacks against GNNs models mainly focus on tasks such as graph classification and node classification, and backdoor attacks against link prediction tasks are rarely studied. In this paper, we propose a backdoor attack against the link prediction tasks based on GNNs and reveal that GNNs are vulnerable to such a threat. The proposed backdoor attack uses a single node as the backdoor trigger and injects backdoor into GNNs models through poisoning selected node pairs in the training graph. In the inference stage, the backdoor will be activated by simply linking the trigger node to the two end nodes of unlinked target node pairs in the input samples and the attacked GNNs models incorrectly predict link relationships between the unlinked target node pairs as linked. We demonstrate that this attack is effective and efficient by performing experimental evaluations on four popular models and four public datasets and comparing them with baseline, and the experimental results show that in the black-box scenario, this attack can achieve high attack success rates (more than 88 %) with small model accuracy losses (less than 1%) and small poisoning rates (about 1%).
Index Terms:
Graph Neural Networks, backdoor attack, link prediction.
I Introduction
@@
Graph-structured data is ubiquitous in the real world, covering various application domains such as social networks [1],urban transportation networks [2], knowledge graphs [3], and biological networks [4]. In order to efficiently process graph-structured data, various powerful graph neural networks(GNNs) have been proposed. GNNs adopt the message passing mechanism to update the representations of nodes by aggregating information from their neighbors. GNNs can be used for various graph related tasks such as graph classification [5], node classification [6], link prediction [7, 8] etc. Link prediction aims to predict the existence of links between pairs of nodes in a graph, which is an important task of GNNs and has a wide range of applications such as friends recommendation in social networks [9], prediction of missing triples in knowledge graphs [10], and prediction of protein functions as well as interactions in biological networks [11].
@@
GNNs have powerful performance in processing graph-structured data. However, recent studies [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22] have shown that GNNs are vulnerable to a new security threat called backdoor attacks. In backdoor attacks, the attacker injects specific patterns (called backdoor triggers) in the training data (the process is called poisoning) and then embeds the backdoor in the GNNs models(called backdoored GNNs models) through training. In the inference phase, when the input samples contain the backdoor triggers, the backdoor in the models will be activated and perform the action specified by the attacker in advance. For example, the GNN models embedded with the backdoor misclassify input samples with triggers as attacker-specified target class label, whereas they perform normally when input samples that do not contain backdoor triggers. Backdoor attacks often occur when the training process of GNN models are uncontrolled, e.g. backdoor attacks can occur by using models trained by third parties or by using third-party data to train the models [22]. Backdoor attacks are highly stealthy and they have posed serious security threat to GNNs.
@@
To the best of our knowledge, the only existed backdoor attack studies on link prediction are LB [20] and DLB [21], where DLB is a backdoor attack method for link prediction on dynamic graphs, and LB is a backdoor attack method for link prediction on static graphs, which studies a similar problem to that of our work. Specifically, LB optimizes the structure and features of a random initial subgraph by using the gradient information of the target model, then obtains a dense subgraph as the trigger for the backdoor attack, and finally poisons the graph dataset with this subgraph, so as to embed the backdoor into the GNN models through training. Although LB achieves good performance, this method requires a large attack budget because LB uses the gradient information of the target model to optimize the subgraph, and the dense subgraph makes the trigger easy to be detected.
@@
In this paper, we investigate whether there are less costly and more stealthy backdoor attacks against link prediction task of GNNs. Specifically, we propose such a backdoor attack with a single node as the trigger, where the backdoored GNNs models incorrectly predict link relationships between two unlinked nodes as linked when the trigger occurs. Initially, a new node is created as the trigger, whose features are generated by counting the frequency of occurrence of each feature of all nodes and selecting the subset of features with fewer occurrences; Then some node pairs in the training graph are selected to inject the trigger for poisoning the training graph, and the backdoor is embedded in the model through the training process. In the inference phase, the backdoor embedded in the model can be activated by simply connecting the trigger node to the two end nodes of the target unlinked node pairs in the test data, causing the backdoored model to incorrectly predict the unlinked two nodes as having a link relationship. The illustration of our backdoor attack is shown in Figure 1. The left side of the figure shows the input graph and the red node is the backdoor trigger. Part (a) of Figure 1 is the input graph without the trigger, and node A and node B is an unlinked target node pair; The input graph in Part(b) is similar to that in Part(a) except that the former contains the trigger connecting to the node A and the node B respectively. When predicting the link relationship between nodes A and B in the input graph without trigger with the backdoored model, it works well and correctly predicts that there is no link between node A and node B, which is represented by 0, while for the input graph with the trigger as shown in Part (b) of Figure 1, the backdoor in the backdoored model is activated due to the presence of the trigger and it predicts incorrectly that there is a link between the node A and the node B, which is represented by 1. We evaluate the performance of the proposed attack on four popular models and four benchmark datasets and compare the results with that of baseline to show that the attack is effective and efficient, and the experimental results indicate that the attack success rate of the attack can reach more than 89 % in the black-box scenario with the poisoning rates of about 1%, and that the classification accuracy of the backdoored model on clean samples is close to that of clean model with the drop less than 1%.
@@
    We propose a backdoor attack against link prediction with GNNs, revealing that GNNs are vulnerable to this security risk. The backdoor attack has following three features: the attack achieves better stealthiness by using only one single node as the trigger, and it requires lower poisoning rate in training dataset ; The attack can be implemented in black box scenarios, where the attackers do not need to know the parameters of target GNN models or create surrogate models, so it is easier and more practical for attacker to launch the attack; The attack has lower computing cost without involving any computation of gradients.
@@
    We propose a simple and efficient method for contaminating training data, which can achieve better attack performance with lower poisoning rates by connecting the trigger node to the selected nodes pairs in the graph.
@@
The rest of the paper is organized as follows. Firstly, we introduce the background knowledge of GNNs, link prediction and backdoor attacks in Section II; Then, we introduce the adversarial attacks against GNNs and further introduce present the state of art of backdoor attacks against GNNs in Section III; Next, we describe in detail our proposed backdoor attack in Section IV and evaluate its performance in Section V;Finally, we conclude the paper in Section VI.
Figure 1: Illustration of our backdoor attack against link prediction. The trigger is the red node and the node A and the node B is an unlinked target node pair. Part (a) is the input graph without the trigger. The backdoored model performs well for the input graph without the trigger and it predicts correctly that there is no link between the node A and the node B, which is represented by 0. The input graph in Part(b) is similar to that in Part(a) except that the former contains the trigger connecting to the node A and the node B respectively. The backdoor in the backdoored model is activated due to the presence of the trigger and it predicts incorrectly that there is a link between the node A and the node B for the input graph with the trigger, which is represented by 1.
II Background
II-A GNNs
Given an undirected and unweighted attribute graph $G=(V,E,X)$, where $V=\{v_{1},v_{2},\cdots,v_{N}\}$ is the set of nodes, $E$ is the set of edges, $X=R^{N\times d}$ represents the nodes feature matrix, and the adjacency matrix $A\in R^{N\times N}$. For two nodes $v_{i},v_{j}\in V$, if $(v_{i},v_{j})\in E$, it means that there exists an edge between $v_{i}$ with $v_{j}$ and $A_{ij}=1$, otherwise, $A_{ij}=0$.$N=|V|$ is the total number of nodes, and $d$ is the number of node feature dimensions.
@@ II-B Link prediction
II-C Link prediction methods based on GNNs
@@
$\begin{split}L(\theta)=&\sum_{(i,j)\in E^{+}}{logp_{\theta}(A_{ij}=1|X)}\\ &+\sum_{(i,j)\in E^{-}}{log(1-p_{\theta}(A_{ij}=1|X))}\end{split}$ (4)
@@
II-D Backdoor attacks
Backdoor attacks are a type of poisoning attack against deep learning models, which aims to implant some hidden malicious functions, i.e., backdoors, in the target models by using poisoned samples with triggers during the models training process. In the inference phase, for input samples that contain trigger, the backdoor will be activated, causing the target models to produce incorrect prediction results, e.g., the input samples are incorrectly misclassified to the target class label specified by the attacker, whereas the backdoored model works normally for benign samples that do not contain trigger. Backdoor attacks were first proposed in the field of images [23, 24]. Gu et al. [23] proposed the backdoor attack for the first time in the image domain, where the backdoored model is trained by poisoning some training samples using special markers as triggers, and in the inference phase, the backdoored model classifies the images with the triggers as the target label, and classifies the clean images normally. Subsequently, backdoor attacks in the textual domain have also been extensively studied [25, 26], e.g., Qi.et.al [26] used syntactic structures as the triggers for textual backdoor attacks.
Backdoor attacks are highly stealthy and harmful because, in the absence of triggers, the performance and behavior of backdoored models are not significantly different from normal models and are not easily detected and defended against. Backdoor attacks can occur in many scenarios, such as when using datasets, platforms, or models provided by third parties for training or deployment. Backdoor attacks pose a serious threat to the security and trustworthiness of deep learning systems, and require sufficient attention and concern.
III Related works
@@ III-A Adversarial attacks
III-B Backdoor attacks
In the graph data field, backdoor attacks have been less studied, and current works focused on backdoor attacks against graph classification [12, 13, 14, 15, 16] and node classification [16, 17, 18, 19]. Xi et al. [12] proposed for the first time a backdoor attack method for GNNs that uses subgraphs as the triggers that can be dynamically customized for different graphs with different triggers to poison the data. Yang et al. [13] revealed a transferable graph backdoor attack without a fixed trigger pattern, which implements black-box attacks on GNN by the attacking surrogate models. Zheng et al [14] rethink the triggers from the perspective of motifs (motifs are frequent and statistically significant subgraphs containing the rich structural information of graphs ), proposed motif-based backdoor attacks, and gave some in-depth explanations of backdoor attacks. Xu et al.[16] select the optimal trigger position through interpretable methods to accomplish backdoor attacks against graph classification and node classification.
The above studies mainly focus on graph classification and node classification, and the only current work on backdoor attacks under the context of link prediction are LB [20] and DLB [21]. DLB investigated the link prediction backdoor attack on dynamic graphs, it generates different initial triggers by Generative Adversarial Network (GAN) [37], and then based on the gradient information of the attack discriminator in the GAN, selects some of the links of the initial triggers to form a trigger set to accomplish the backdoor attack. LB is similar to our work, but LB uses subgraphs as triggers and generates triggers by using the gradient information of the target models. Compared to LB, Our proposed backdoor attack has following three different features: the attack achieves better stealthiness by using only one single node as the trigger, and it requires lower poisoning rate in training dataset; The attack can be implemented in black box scenarios, where the attackers do not need to know the parameters of target GNN models or create surrogate models, so it is easier and more practical for attacker to launch the attack; The attack has lower computing cost without involving any computation of gradients.
IV Method
In this section, we explain in detail how our attack is implemented. Table I summarizes the notions used in the following sections and their explanations.
Table I: The explanations of the notions
Notations Explanations
$G=(V,E,X)$ Graph $G$ with edge set $E$, node set $V$ and feature matrix $X$
@@ IV Method
$f_{c}$ The clean GNN model for link prediction
@@
$x_{t}$ The features vector of the trigger node
$f_{b}$ The backdoored GNN model for link prediction
$T$ The linked state
$E_{ij}$ The target unlinked node pair without the trigger
$\widetilde{E_{ij}}$ The target unlinked node pair with the trigger
$\alpha_{j}$ The frequency of occurrence of ”1” of the j-th feature
$M_{k}$ The top k indexes of dimension where the nodes with the least frequent occurrence
@@
$Score$ The poisoned node pair scoring function
IV-A Attack overview
The backdoor attack proposed in this paper aims to the link prediction task under transductive learning, the goal of which is to incorrectly predict a link relationship between two unlinked target nodes. Formally, given a graph $G$ and a trigger $n_{t}$, the attacker generates the poisoned graph $\tilde{G}$ by embedding the trigger $n_{t}$ into the graph $G$. The backdoored model $f_{b}$ is obtained by training on $\tilde{G}$ and the clean model can be denoted as $f_{c}$. The backdoor attack target can be denoted as: